OpenStudy (anonymous):

Requesting a double check of my answers and a step by step for question "b" please. Answer the questions below about the quadratic function. f(x)=-3x^2+6x-6 a. Does the function have a minimum or maximum value? Answer: Maximum- the parabolas vertex is shifted downward b. Where is the functions minimum or maximum value? Answer: ????? c. Where does the minimum or maximum value occur? Answer: x= (1,-3)

OpenStudy (anonymous):

a) answer : maximum, because the leading coefficient it negative, therefore the parabola opens down. not because of the vertex

OpenStudy (anonymous):

b) the function has a maximum value at the second coordinate of the vertex the first coordinate of the vertex is \(-\frac{b}{2a}=-\frac{6}{2\times (-3)}=1\) and the second coordinate of the vertex is \(-3\times 1^2+6\times 1-6=-3+6-6=-3\)

OpenStudy (anonymous):

the short answer to b) is \(-3\) and the short answer to c) is \(1\)

OpenStudy (anonymous):

For a. because (a<0) and the leading coefficient is a negative number- gotcha!

OpenStudy (anonymous):

i am a little confused on the wording to b) i take it to mean WHAT is the functions maximum value

OpenStudy (anonymous):

"where" is a bit ambiguous

OpenStudy (anonymous):

yes for "b" it is what is the functions maximum value.

OpenStudy (anonymous):

ok, in that case the answer is "the maximum value is \(-3\)"

OpenStudy (anonymous):

because the extreme value is -3 right?

OpenStudy (anonymous):

for b. I'm still confused- it looks like 1 is on the x axis and -3 is on the y. I won't have to put both down for b? Just the y axis?

OpenStudy (anonymous):

so it wouldn't be (1,-3) it would just be -3

OpenStudy (anonymous):

Thanks again satellite!